# CLIP-Free, Label-Free, Zero-Shot Concept Bottleneck Models

### v3

@@ Appendix B Limitations
As with any research work, our study has its own limitations that should be transparent and acknowledged. In particular, we identify a primary limitation of our method concerning wrong semantic associations of class names in CBMs. This issue is closely tied to the choice of concept set used. When using the 20K most common words in English as our concept set, we observe that class names get associated to wrong semantically-related concepts. Figure 1 illustrates such cases. In the first example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric). This occurs because the bird “drake" is less familiar to the text encoder than the artist “drake". In fact, a google search with the word “drake” yields directly the artist rather than the bird. In the second example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself. We also observe similar cases that may raise ethical concerns (e.g., the animal “cock" leads to associations with male reproductive terms). For each example, we also report the total logit score of the prediction.
However, it is worth noting the following:
@@
Figure 1: Limitations of our method in wrong semantic concept association
Appendix C Ablation Studies of the MLP
C.1 Role and Impact of the MLP
@@ C.2 MLP Design
Appendix D Ablation Studies on the Text Encoder
In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library (reimers-2019-sentence-bert) with the ResNet50 visual classifier. We observe that the choice of the text encoder has minimal effect on the performance. This is because even lower-performing text encoders are capable of understanding class names.
Text Encoder Top-1 (%)
DistilRoberta 75.73
MPNet-Base 75.78
MPNet-Base-MultiQA 75.76
MiniLM 75.80
Table 3: Ablation studies on other text encoders
Appendix E Effectiveness of Concept Interventions
@@
As a supplementary experiment, we also report concept intervention results on our ZS-CBMs. Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions. The Waterbirds-100 dataset (Sagawa2020Distributionally) is a standard dataset used in previous works (Rao2024DiscoverthenNameTC) to conduct CBM intervention experiments. It is a binary classification dataset of two classes: waterbirds and landbirds. The training images of waterbirds are on water backgrounds, and training images of landbirds are on land backgrounds. However, the validation images do not have that correlation, where waterbird images appear on land backgrounds and landbird images on water backgrounds. The model is assumed to learn the water–land background correlation to perform this classification task. By building a CBM, we can correct this bias by intervening in concepts in the CB layer. We curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set. Our curated dataset includes 140 validation images (70 for each class). We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from (Rao2024DiscoverthenNameTC) which includes a collection of bird-related concepts and a collection of land-related concepts. The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification. To correct this, we intervene in the concepts in the CB layer, following the setup from (Rao2024DiscoverthenNameTC). Intervention R (Int. R): We zero-out activations of any bird concepts from the bottleneck layer, and expect the accuracy to drop. Intervention K (Int. K): We keep activations of bird concepts as they are, but scale down the activations of all remaining concepts by multiplying them with a factor of 0.1, and expect the accuracy to increase. Results are presented in Table 4 for some models, and demonstrate the success of our intervention experiments.
@@
We also conduct concept intervention experiments with a more challenging multi-class setup. Unlike the Waterbirds dataset where bias‑correlation issues are assumed, we make no such assumption here. We instead evaluate whether intervening on class-related concepts in the bottleneck layer and zeroing out their activations, impairs model accuracy. A drop in accuracy indicates that these concepts are important for prediction. We select a subset of 10 classes from ImageNet following (Howard_Imagenette_2019). We use the original ImageNet validation images for those classes, rather than the validation set from (Howard_Imagenette_2019), because the latter contains original ImageNet training examples that our model has already seen. We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts. The classes and concepts we used are provided in Section Q of the supplementary material. We then measure CBM accuracy before and after intervention. Results are provided in Table 5. For each image, we zero out the activations of its class-related concepts and report the Intervention R (Int. R) metric. As shown in Table 5, this intervention reduces accuracy by approximately 20% on average, underscoring the concepts importance to the CBM.
@@
Appendix F Transformed Classifier Results on Other Datasets
@@
Appendix G CBM Results on Other Datasets
We report CBM results on additional datasets: Places365 (domain-specific to scenes), DTD (domain-specific to texture and fine-grained), and EuroSAT (domain-specific to satellite images). When a baseline method is also reported on that dataset, we include it as a baseline. Otherwise, we use CLIP models as a baseline to act as a feature extractor and to compute concept activations, and train a linear classifier on top of the concept activations, formulating a CLIP supervised baseline. All baselines use the same concept set. Results are shown in Table 7. For Places 365, we can see that the transformed DenseNet161 classifier trained only on ImageNet, outperforms supervised CLIP-based ResNet and ViTs CBM methods in a zero-shot manner. The same applies for EuroSAT and DTD with our transformed Image-Net only trained classifiers. Therefore, the experiments show that our method scales to domain-specific and fine-grained datasets.
Dataset Method Model Acc (%)
Places365 Supervised
DCLIP CLIP-ResNet50 37.90
DCLIP CLIP-ViT-B/16 40.30
LF-CBM CLIP-ResNet50 49.00
LF-CBM CLIP-ViT-B/16 50.60
CDM CLIP-ResNet50 52.70
CDM CLIP-ViT-B/16 52.60
Zero-Shot
Ours ResNet50 51.57
Ours DenseNet161 53.42
EuroSAT Supervised
Baseline CLIP-ResNet50 86.27
Baseline CLIP-ViT-B/16 88.57
Zero-Shot
Ours ResNet50 94.22
Ours WideResNet101 94.12
Ours ViT-B/16 93.65
DTD Supervised
Baseline CLIP-ResNet50 57.77
Baseline CLIP-ViT-B/16 61.86
Zero-Shot
Ours ResNet50 68.88
Ours WideResNet101 66.97
Ours ViT-B/16 68.46
Table 7: CBM results on Places365, EuroSAT, and DTD datasets
Appendix H Prompt Variations
@@
Appendix I Other Concept Sets
Early works such as label-free CBMs use LLMs to automatically generate a set of concepts for a target class. In recent works such as DN-CBM (Rao2024DiscoverthenNameTC) and DCBM (Knab2024DCBMDV), it was however shown that general, pre-defined concept sets outperform the LLM-generated ones, which introduce many spurious correlations and hallucinated associations. In Table 9, we compare general pre-defined and LLM-generated concept sets. Across different types of architectures, general, pre-defined concept sets outperform the LLM-generated ones by several percent. Please note that while earlier works such as LF-CBMs employ LLM-generated concept sets (which underperform compared to the general, pre-defined concept set), all results reported in our main manuscript—including those from prior works—are presented using the general, pre-defined concept set to ensure a fair comparison.
@@ Appendix I Other Concept Sets
Appendix J Implementation Details
For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library (reimers-2019-sentence-bert). This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs. It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization (Ba2016LayerN), a GELU activation function (Hendrycks2016GaussianEL) and Dropout (Srivastava2014DropoutAS) with a drop probability of 0.5. The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function. The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder). We train the MLP projector with a batch size of 256 using the ADAM optimizer (KingBa15) with a learning rate of 1e-4 that decays using a cosine schedule (loshchilov2017sgdr) over the total number of epochs. We follow the original image sizes that the classifier was trained on.
@@
Appendix K Compositional Image Captioning
Compositional Image Captioning is an old image captioning paradigm termed (Kulkarni2011BabyTU), later revived with deep learning methods (Lu2018NeuralBT). In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs) and their powerful capabilities, we use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 3.2 and shown in Figure 2(a), and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 3.2. We also add a list of the most common verbs (commonverbs) in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a limited set of in-context examples from the COCO captioning training set (in our experiments, we use 9 examples). This allows us to generate sentences adhering to a specific style and structure. For this experiment, we used GPT4o-mini (gpt4omini) as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importance. We use the following prompt:
@@
K.1 Domain-Specific Compositional Captioning
Since the LLM we used for compositional image captioning is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains. We explore alternative concept sets in Compositional Captioning. In the main manuscript, we reported results using the 20,000 most common English words as our concept set. Since the LLM remains fixed and functions as a composer, integrating detected concepts and verbs grounded in the image into a caption, we can seamlessly substitute the concept set with any domain-specific concept set alternative. This allows for the generation of captions tailored to a specific domain. Here, we maintain the same set of verbs but explore the use of concepts specific to the ImageNet dataset. Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset. We use the ImageNet-specific concept set from (oikarinen2023labelfree) and report zero-shot captioning performance in Table 11. As shown, we observe a decrease in all metrics. This shows that our method can readily produce captions for any domain. Finally, also note that we can control the style of the generations by simply prompting the LLM to compose the concepts and verbs in a specific style (e.g., humorous, positive, negative).
@@
Appendix L Performance on Additional Models
@@
Appendix M Process of Zero-Shot Image Captioning
We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$. A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$. To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning (Li2021PrefixTuningOC), which prepends learnable tokens in the embedding space. We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly. Our method builds upon the work of (Tewel2021ZeroCapZI).
@@
We also add the fluency loss from (Tewel2021ZeroCapZI) as well as other token processing operations. We refer readers to (Tewel2021ZeroCapZI) for more information. We use the small GPT-2 of 124M parameters as $G$. We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
@@
Appendix O Using only class names
We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP. In practice, we can go beyond class names by using resources like a class hierarchy from WordNet (Lin1998WordNetAE) (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL (Pratt2022WhatDA) or VCVD(Menon2022VisualCV). However, this approach would be considered as “cheating. The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes. Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn. Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation. For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts. For these reasons, we refrain from using any other additional information than the class names. We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
Appendix P Global Class-Wise Qualitative Examples
In Figure 3, we present a probability distribution of global class-wise concepts. These are concepts detected for all images of a specific class, along with their frequency. We consider two semantically similar classes but distinctively different: “hammerhead shark" and a “tiger shark". We highlight in yellow the top concepts in “hammerhead shark" that are not present in “tiger shark". These concepts are “harpoon" and “lobster hammer", both which are distinctive to the head of the hammerhead shark and drive its prediction.

### v4

@@ 1 Limitations
As with any research work, our study has its own limitations that should be transparent and acknowledged. In particular, we identify a primary limitation of our method concerning wrong semantic associations of class names in CBMs, due to polysemy. This issue is closely tied to the choice of concept set used. When using the 20K most common words in English as our concept set, we observe that class names get associated to wrong semantically-related concepts. For example, the bird “drake" is linked to artist-related concepts (Rihanna, Robbie, lyric). This occurs because the bird “drake" is less familiar to the text encoder than the artist “drake". In fact, a google search with the word “drake” yields directly the artist rather than the bird. As another example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself. We also observe similar cases that may raise ethical concerns (e.g., the animal “cock" leads to associations with male reproductive terms). However, it is worth noting the following:
@@
2 Ablation Studies on the Text Encoder
In Table 1, we present ablation studies using other text encoders from the Sentence-BERT library [45] with the ResNet50 visual classifier. We observe that the choice of the text encoder has minimal effect on the performance. This is because even lower-performing text encoders are capable of understanding class names.
Text Encoder Top-1 (%)
DistilRoberta 75.73
MPNet-Base 75.78
MPNet-Base-MultiQA 75.76
MiniLM 75.80
Table 1: Ablation studies on other text encoders
3 Ablation Studies of the MLP
3.1 Role and Impact of the MLP
@@ 3.2 MLP Design
4 Concept Interventions and other Explainability Metrics
4.1 Concept Interventions
@@
We report concept intervention results on our ZS-CBMs. Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions. The Waterbirds-100 dataset [46] is a standard dataset used in previous works [44] to conduct CBM intervention experiments. It is a binary classification dataset of two classes: waterbirds and landbirds. The training images of waterbirds are on water backgrounds, and training images of landbirds are on land backgrounds. However, the validation images do not have that correlation, where waterbird images appear on land backgrounds and landbird images on water backgrounds. The model is assumed to learn the water–land background correlation to perform this classification task. By building a CBM, we can correct this bias by intervening in concepts in the CB layer. We curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set, including 140 validation images (70 for each class). We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from [44] which includes a collection of bird-related concepts and a collection of land-related concepts. The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification. To correct this, we intervene in the concepts in the CB layer, following the setup from [44]. Intervention R (Int. R): We zero-out activations of any bird concepts from the bottleneck layer, and expect the accuracy to drop. The larger the drop, the better. Intervention K (Int. K): We keep activations of bird concepts as they are, but scale down the activations of all remaining concepts by multiplying them with a factor of 0.1, and expect the accuracy to increase. More increase is better. Results are presented in Table 4 for some models, and demonstrate the success of our intervention experiments.
@@
We also conduct concept intervention experiments with a more challenging multi-class setup. Unlike the Waterbirds dataset where bias‑correlation issues are assumed, we make no such assumption here. We instead evaluate whether intervening on class-related concepts in the bottleneck layer and zeroing out their activations, impairs model accuracy. A drop in accuracy indicates that these concepts are important for prediction. We select a subset of 10 classes from ImageNet following [16]. We use the original ImageNet validation images for those classes, rather than the validation set from [16], because the latter contains original ImageNet training examples that our model has already seen. We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts. The classes and concepts we used are provided in Section 16 of the supplementary material. We then measure CBM accuracy before and after intervention. Results are provided in Table 5. For each image, we zero out the activations of its class-related concepts and report the Intervention R (Int. R) metric. As shown in Table 5, this intervention reduces accuracy by approximately 20% on average, underscoring the concepts importance to the CBM.
@@
4.2 Concept Discovery Accuracy
We use the discrete-based Mutual Information (MI) metric [^2] https://github.com/fawazsammani/clip-interpret-mutual-knowledge to calculate the MI between concepts detected by our method, and those detected by CLIP. Specifically, given the concept bank, we use the CLIP Vision Encoder to encode an image, and the CLIP Text Encoder to encode the concept bank. We record the set of top-K detected concepts most similar to the image. We then encode the same image with our TextUnlock-ed visual classifier (e.g., DINOv2)(including the MLP) and encode the same concept bank with our Text Encoder $T$. We record the set of top-K detected concepts most similar to the image. We then calculate the MI between the detected concepts from CLIP (CLIP ViT-B/16 model) and our method. Results are shown in Table 6 and demonstrate high mutual information which indicates that our method successfully detects key concepts in the dataset by only learning from the limited vocabulary of class labels.
Model NMI
DINOv2 0.681
EfficientNetv2-M 0.704
ViT-B/16v2 0.711
Table 6: Normalized Mutual Information (NMI) with CLIP ViT-B/16 for different models. NMI ranges from 0 (no mutual information) to 1 (perfect agreement).
5 Transformed Classifier Results on Other Datasets
@@
6 Prompt Variations
@@
7 Other Concept Sets
Early works such as label-free CBMs use LLMs to automatically generate a set of concepts for a target class. In recent works such as DN-CBM [44] and DCBM [19], it was however shown that general, pre-defined concept sets outperform the LLM-generated ones, which introduce many spurious correlations and hallucinated associations. In Table 9, we compare general pre-defined and LLM-generated concept sets. Across different types of architectures, general, pre-defined concept sets outperform the LLM-generated ones by several percent. Please note that while earlier works such as LF-CBMs employ LLM-generated concept sets (which underperform compared to the general, pre-defined concept set), all results reported in our main manuscript—including those from prior works—are presented using the general, pre-defined concept set to ensure a fair comparison.
@@ 7 Other Concept Sets
8 Implementation Details
For the text encoder, we use the all-MiniLM-L12-v1333https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library [45]. This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs. It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization [2], a GELU activation function [15] and Dropout [49] with a drop probability of 0.5. The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function. The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder). We train the MLP projector with a batch size of 256 using the ADAM optimizer [18] with a learning rate of 1e-4 that decays using a cosine schedule [29] over the total number of epochs. We follow the original image sizes that the classifier was trained on.
@@ 8 Implementation Details
9 Compositional Image Captioning
Compositional Image Captioning is an old image captioning paradigm termed [21], later revived with deep learning methods [30]. In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs) and their powerful capabilities, we use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 3.2 and shown in Figure 2(a), and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 3.2. We also add a list of the most common verbs [8] in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a limited set of in-context examples from the COCO captioning training set (in our experiments, we use 6 examples). This allows us to generate sentences adhering to a specific style and structure of the in-context examples. For this experiment, we used GPT4o-mini [37] as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importance. We use the following prompt:
@@ 9 Compositional Image Captioning
9.1 Domain-Specific Compositional Captioning
Since the LLM we used for compositional image captioning is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains. We explore alternative concept sets in Compositional Captioning. In the main manuscript, we reported results using the 20,000 most common English words as our concept set. Since the LLM remains fixed and functions as a composer, integrating detected concepts and verbs grounded in the image into a caption, we can seamlessly substitute the concept set with any domain-specific concept set alternative. This allows for the generation of captions tailored to a specific domain. Here, we maintain the same set of verbs but explore the use of concepts specific to the ImageNet dataset. Since ImageNet lacks dedicated captions, we evaluate the domain-specific captioning by anticipating a decline in performance on the COCO captioning dataset. We use the ImageNet-specific concept set from [36] and report zero-shot captioning performance in Table 11. As shown, we observe a decrease in all metrics. This shows that our method can readily produce captions for any domain. Finally, also note that we can control the style of the generations by simply prompting the LLM to compose the concepts and verbs in a specific style (e.g., humorous, positive, negative).
@@ 9.1 Domain-Specific Compositional Captioning
10 Qualitative examples of zero-shot image captioning.
We present qualitative examples of zero-shot image captioning from different classifiers in Figure 1. We choose BeiT-L/16, ConvNeXtv2, and ViT-B/16. This allows us to see how different classifiers “see" the image. From the first example, BeiT-L/16 captures features of both the vegetables and the dog, whereas ConvNeXtV2 captures only the vegetables. In contrast, ViT-B/16 focuses exclusively on the dog and its characteristics.
Figure 1: Qualitative examples of zero-shot image captioning.
11 Performance on Additional Models
@@
[…]
BeiT-B/16 84.54 85.06 -0.52
WideResnet50 78.35 78.47 -0.12
[…]
EfficientNetv2-S 84.04 84.23 -0.19
ConvNeXt-Small 83.42 83.62 -0.20
ResNeXt101-32x8d 79.10 79.31 -0.21
[…]
DenseNet169 75.46 75.60 -0.14
@@
12 Process of Zero-Shot Image Captioning
We remind readers of the mapping function, denoted as MLP, that transforms the visual features $f$ into the same space as textual features, producing $\tilde{f}$. A pre-trained language model $G$ is then optimized to generate a sentence that closely aligns with $\tilde{f}$. To preserve the generative power of $G$, we keep it frozen and apply prefix-tuning [22], which prepends learnable tokens in the embedding space. We follow a test-time training approach to optimize learnable tokens for each test input on-the-fly. Our method builds upon the work of [51].
@@ 12 Process of Zero-Shot Image Captioning
We also add the fluency loss from [51] as well as other token processing operations. We refer readers to [51] for more information. We use the smallest GPT-2 of 124M parameters as $G$. We also noticed that using a bigger $G$ (e.g., GPT-2 medium) does not enhance performance, indicating that a decoder with basic language generation knowledge is sufficient.
@@ 13 Concept Filtering
14 Using only class names
We remind readers from Section 3 that we only use the class names to formate the text prompt for the text encoder when training the MLP. In practice, we can go beyond class names by using resources like a class hierarchy from WordNet [24] (the original source where ImageNet was extracted from), or class descriptions extracted from a LLM as in CuPL [42] or VCVD[33]. However, this approach would be considered as “cheating. The original classifier implicitly learns the semantics, hierarchies, relationships and distinctive features of different classes. Explicitly providing additional information would not replicate the classifier faithfully since it would force the classifier to focus on predefined features or those we intend it to learn. Moreover, this would also leak information to downstream tasks such as CBMs and textual decoding of visual features, compromising the fairness of evaluation. For instance, if class descriptions were used in the training, the concepts in CBMs would align with those specified in the training prompts. For these reasons, we refrain from using any other additional information than the class names. We use the class names provided from https://gist.github.com/yrevar/942d3a0ac09ec9e5eb3a.
15 Global Class-Wise Qualitative Examples
In Figure 3, we present a probability distribution of global class-wise concepts. These are concepts detected for all images of a specific class, along with their frequency. We consider two semantically similar classes but distinctively different: “hammerhead shark" and a “tiger shark". We highlight in yellow the top concepts in “hammerhead shark" that are not present in “tiger shark". These concepts are “harpoon" and “lobster hammer", both which are distinctive to the head of the hammerhead shark and drive its prediction. Note that for this experiment, we do not apply the concept filtering procedure.